OpenStudy (anonymous):

Medal award!! ^^ Idk how to write a part of the question, so i'll draw the whole thing out ^^

OpenStudy (anonymous):

 (A) 3 (B) 9/2 (C) 9 (D) 18 (E) 81 Is it A?

OpenStudy (anonymous):

\[r=\sqrt{3v/\pi^h}\]Let k be the number we are looking for:\[9r=\sqrt{3kv/\pi^h}\]Divide second expression by the first one:\[9=\sqrt{k}\rightarrow k=81\]

OpenStudy (anonymous):

Ah...Thanks!